(a) What is the expression for the magnitude of the force F required to accelerate an elevator of mass m upward starting from rest to a height H to a final speed of V(f)?

(b) What is the value of the force if m=1800 kg, h=120 m, and v(f)=16m/s?
Answer #1

a)

let the acceleration of the elevator is a

using third equation of mtoion

V(f)^2 = 2 * H * a

a = V(f)^2/(2 *H)

force required = m * a

force required = m * V(f)^2/(2 *H)

b)

for the value of F

F = 1800 * 16^2/(2 * 120)

F = 1920 N

the value of force F is 1920 N